\begin{thebibliography}{99}
\vspace{-10pt}
\bibitem{Aur99} \joref{Aurich R.}{\apj}{524}{497}{1999} 
\ (arXiv:astro-ph/9903032)
\bibitem{BPS98}  
  \joref{Bond J.~R., Pogosyan D., 
Souradeep T.}{\cqg}{15}{2573}{1998}\ (arXiv:astro-ph/9804041); 
arXiv:astro-ph/9912124, arXiv:astro-ph/9912144 
%
%\bibitem{BPS00a}  
%  \epref{Bond J.~R., Pogosyan D., 
%Souradeep T.}{arXiv:astro-ph/9912124}{2000a}
%\bibitem{BPS00b}  
%  \epref{Bond J.~R., Pogosyan D., 
%Souradeep T.}{arXiv:astro-ph/9912144}{2000b}

\bibitem{CS99} 
\epref{Cornish N.~J., Spergel D.~N.}{arXiv:astro-ph/9906401}{1999}


\bibitem{Corn96} Cornish N.~J., Spergel D.~N., Starkman G.~D., 1996, 
arXiv:gr-qc/9602039

\bibitem{Corn98b} \joref{Cornish N.~J., Spergel D.~N., 
Starkman G.~D.}{\cqg}{15}{2657}{1998}\ (arXiv:astro-ph/9801212) 
%% circles method


\bibitem{FagG97} \joref{Fagundes H. \& 
Gausmann E.}{PhysLettA}{238}{235}{1998} 
\ (arXiv:astro-ph/9704259) %% Euclidean spikes

\bibitem{FagG99a} Fagundes H. \& Gausmann E., 1999a, in 
Proceedings of the XIX Texas Symp. Rel. Astr. Cosmology, 
CD-ROM version (arXiv:astro-ph/9811368) %% CC in CHMs

\bibitem{G99a} \epref{Gomero G.~I., Teixeira A.~F.~F., 
Rebou\c{c}as M.~J., Bernui A.}{arXiv:gr-qc/9811038}{1999a} %% paper I 
\bibitem{G99b} \joref{Gomero G.~I., Rebou\c{c}as M.~J., 
Teixeira A.~F.~F.}{PhysLettA}{in press}{}{2000} 
\ (arXiv:gr-qc/9909078)  %% cas euclidien ``paper II''
\bibitem{G99c} \epref{Gomero G.~I., Rebou\c{c}as M.~J., 
Teixeira A.~F.~F.}{arXiv:gr-qc/9911049}{1999c} %% signatures topologiques

\bibitem{Gott80} \joref{Gott, J.R.I.}{MNRAS}{193}{153}{1980}


\bibitem{Inoue99} 
\epref{Inoue K.~T.}{arXiv:astro-ph/9810034}{1999}

\bibitem{LaLu95} \joref{\mbox{Lachi\`eze--Rey} M.,  
Luminet J.--P.}{Phys. Rep.}{254}{136}{1995} 
\ (arXiv:gr-qc/9605010)

\bibitem{LevSGSB98} \epref{Levin J.,
Scannapieco E., de Gasperis, G., Silk J., Barrow,
J.D.}{arXiv:astro-ph/9807206}{1999} %% How the Univ got its spots


\bibitem{LLL96} \joref{Lehoucq R., Luminet J.--P., 
Lachi\`eze--Rey M.}{\aanda}{313}{339}{1996} \ (arXiv:gr-qc/9604050)
\bibitem{LLU98} \joref{Lehoucq R., Luminet J.--P., 
Uzan J.--Ph.}{\aanda}{344}{735}{1999} \ (arXiv:astro-ph/9811107)

\bibitem{LR99} Luminet J.--P., Roukema B.~F., 1999, 
`Topology of the Universe: Theory and Observations', 
Carg\`ese summer school `Theoretical and Observational Cosmology', ed. 
Lachi\`eze-Rey M., Netherlands: Kluwer, p117 \ (arXiv:astro-ph/9901364)

\bibitem{Rouk96} \joref{Roukema B.~F.}{\mnras}{283}{1147}{1996} 
%\ (arXiv:astro-ph/9603052)

\bibitem{Rouk00a} \joref{Roukema B.~F.}{\mnras}{312}{712}{2000a} 
\ (arXiv:astro-ph/9910272)
\bibitem{Rouk00b} \joref{Roukema B.~F.}{submitted}{}{}{2000b} 
{\em Large Scale Structure Constraints on Cosmic Topology: 
Nested Crystallography?}
\bibitem{Rouk00c} \joref{Roukema B.~F.}{\cqg}{17}{3951}{2000c} 
\ (arXiv:astro-ph/0007140)
%{\em A Counterexample to Claimed COBE Constraints on Compact Toroidal Models}
\bibitem{RE97} \joref{Roukema B.~F., Edge A.~C.}{\mnras}{292}{105}{1997} 
\ (arXiv:astro-ph/9706166) 

\bibitem{UzPet97} \joref{Uzan J.--Ph., 
Peter P.}{Phys. Lett.}{B406}{20}{1997}

\bibitem{ULL99} \joref{Uzan J.--Ph., Lehoucq R. \& Luminet J--P.}
{\aanda}{351}{766}{1999} \ (arXiv:astro-ph/9903155)

\bibitem{LUL00} \joref{Lehoucq R., Uzan J.--Ph.  \& Luminet J--P.}
{submitted}{}{}{2000} \ (arXiv:astro-ph/0005515)

%\bibitem{Wein72} Weinberg S., 1972, Gravitation 
%and Cosmology, New York, U.S.A.: Wiley
\bibitem{Wich99} Wichoski, U., 1999, in 
CTP98 proceedings, arXiv:astro-ph/0010170
%{\it Cosmological Topology in Paris 1998} Workshop, 
%(Obs. Paris/IAP, Paris, 14 Dec 1998), %\\
%{\em http://www.iap.fr/users/roukema/CTP98/programme.html}
%\vspace{-10pt}
\end{thebibliography}

\documentclass{ws-p9-75x6-50}

\begin{document}

\title{Observational Approaches to the Topology of the Universe}

\author{B. F. Roukema}

\address{Inter-University Centre for Astronomy and Astrophysics, 
 Post Bag 4, Ganeshkhind, Pune, 411 007, India\\
 Email: boud@iucaa.ernet.in}


\maketitle

\abstracts{
Many different and complementary strategies
for translating the basic principle of multiple topological imaging into
observational analysis are now available,
both for three-dimensional and two-dimensional catalogues.}

\section{Introduction}
Observational cosmic topology\cite{LaLu95,LR99}
shares the fundamental observational
problems of curvature estimates (evolution of objects, 
peculiar velocities, etc., see Section 5.3 of ref.\cite{LR99}), 
but since the former concerns global
geometry and the latter only local geometry, the former requires additional 
work in classifying the methods from a purely geometrical point
of view. Most work is based on the principle of 
the existence of multiple topological images of single physical objects.


\section{Complementary geometrical strategies}

{
\begin{list}{\Alph{enumi}.}{\usecounter{enumi}} \parskip=-8pt
\item multiple topological images:
  \begin{list}{\Alph{enumi}.\roman{enumii}}{\usecounter{enumii}}
     \item three-dimensional (collapsed astrophysical objects):
  \begin{list}
{\Alph{enumi}.\roman{enumii}.\arabic{enumiii}}{\usecounter{enumiii}}
     \item local isometries\cite{Rouk96,ULL99,LUL00} - multiple occurrence 
of ``type I pairs'' or ``local pairs''
     \item cosmic 
crystallography\cite{LLL96,LLU98,G99a,FagG97,FagG99a,G99b,G99c,ULL99,LUL00}
 - multiple occurrence of ``type II pairs'' or ``generator pairs'',
i.e. of pairs of objects in comoving space separated by a generator
     \item[$\bullet$] for comparison, the uncorrelated pairs
from simply connected random 
simulations could be called ``type III pairs'' or ``random pairs''
     \item characteristics of individual objects\cite{Gott80,RE97,Wich99}
     \end{list}

     \item two-dimensional (microwave background, CMB):
     \begin{list}
{\Alph{enumi}.\roman{enumii}.\arabic{enumiii}}{\usecounter{enumiii}}
	\item identified circles principle: discovery of 
principle\cite{Corn96,Corn98b} and its quantitative application to COBE 
data\cite{Rouk00a,Rouk00c}
	\item patterns of spots\cite{LevSGSB98}
	\item perturbation statistics assumptions: see refs listed under 
(i) in Section 1.2 of ref.\cite{Rouk00a}; 
controversy 
against\cite{BPS98} and
in favour of\cite{Inoue99,Aur99,CS99} 
hyperbolic compact 3-manifolds 
by applying these assumptions to COBE data remains
     \end{list}
  \end{list}
\item other:
  \begin{list}{\Alph{enumi}.\roman{enumii}}{\usecounter{enumii}}
    \item cosmic strings\cite{UzPet97}
    \item nested crystallography\cite{Rouk00b}
  \end{list}
\end{list}
}  



\def\apj{ApJ.}                        
\def\apjs{ApJ. Supp.}                 
\def\aj{A.J.}                           
\def\aanda{A\&A}                      
\def\cqg{Class. Quant. Grav.}         
\def\mnras{Month. Not. R. Astron. Soc.}

\newcommand\joref[5]{#1, #5, {#2, }{#3, } #4}
\newcommand\epref[3]{#1, #3, {#2}}


\end{document}